OpenStudy (anonymous):

Given the linear equation y=1/2x-5, find the y-coordinates of the points (-8, ), (-2, ), and (4, )

OpenStudy (anonymous):

substitute x = -8 , -2 and 4 and find the corresponding values of y

OpenStudy (anonymous):

is this like this, y=F(x)=1/(2x-5) F(-8)=1/(2(-8)-5)=-1/21 F(-8)=1/(2(-2)-5)=-1/9 F(-8)=1/(2(4)-5)=1/3

OpenStudy (anonymous):

good luck bella